OpenStudy (anonymous):

Jake wrote this equation for the area of a rectangle. 431.2 = 24.5w. What is the width (w) of this rectangle? A. 406.7 units B. 67.7 units C. 17.6 units D. 10,564.4 units

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Take 431.2 and divide by 24.5

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

17.6 units

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

to find width, it would be very easy to have an expression defined by only w this can be done by either dividing both sides by 24.5 or subtracting 23.5w from both sides subtracting both sides by 23.5w will yield w = 431.2 -23.5w this doesn't really help so let's divide both sides by 24.5 24.5w/24.5 = 431.2/24.5 w = 17.6

OpenStudy (anonymous):

can you help me with more

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

The length of one side of a square is cm. What is the area of the square? A. 43 cm2 B. 172 cm2 C. 86 cm2 D. 7396 cm2

OpenStudy (anonymous):

i think you left out some vital information

OpenStudy (anonymous):

opps wait